OpenStudy (katherinesmith):

what is the 20th term of the following arithmetic sequence?

OpenStudy (katherinesmith):

\[\frac{ 1 }{ 2 }, 1, \frac{ 3 }{ 2 }, 2\]

OpenStudy (katherinesmith):

@zepdrix

zepdrix (zepdrix):

\[\Large a_1=\frac{1}{2}, \qquad a_2=1\] If we subtract,\[\Large a_2-a_1=1-\frac{1}{2}\]Let's call this value k.\[\Large k=\frac{1}{2}\] This is the common difference between each term.

zepdrix (zepdrix):

The nth term is given by,\[\Large a_n=a_1+(n-1)k\]

zepdrix (zepdrix):

\[\Large a_{20}=?\]Can you figure this one out? :) Plug in 20 for all of your n's. And plug in the first value for a_1.

OpenStudy (katherinesmith):

a20 = 30?

zepdrix (zepdrix):

Hmm let's fill in the pieces.\[\Large a_\color{royalblue}{n}=\color{orangered}{a_1}+(\color{royalblue}{n}-1)\color{green}{k}\] \[\Large \color{royalblue}{n=20},\qquad\qquad \color{orangered}{a_1=\frac{1}{2}},\qquad\qquad \color{green}{k=\frac{1}{2}}\]

zepdrix (zepdrix):

\[\Large a_\color{royalblue}{20}=\color{orangered}{\frac{1}{2}}+(\color{royalblue}{20}-1)\color{green}{\frac{1}{2}}\]

zepdrix (zepdrix):

Try to solve that! :D